OpenStudy (anonymous):

if f(x)=3x^2 - 7x, find the equation of the tangent line at x=2

OpenStudy (anonymous):

The tangent line at x = 2 is \(y-y_0=m(x-x_0),\text{ where }(x_0,y_0)=(2,f(2)) \text{ and }m=f'(2).\)